. Calculate the temperature at which a solution containing 54 g of glucose, (C6H12O6), in 250 g of water will freeze.
(Kf for water = 1.86 K mol-1 kg)

Grade:

1 Answers

Sunil Kumar FP
9 years ago
molality of the glucose =no of mole of glucose in 1 kg of water
no of mole of glucose =weight of the glucse/molecular weight of the glucose =54/180 =.3
we have ,delta Tf=Kfm
delta Tf= (1.86*.3)/.250
=2.232
temperature at which it will freeze=273 -2.232 K =270.768 K